OpenStudy (anonymous):

f(x)=x^2-8x-5 what is the y-coordinate of the vertex? the minimum? what is the range of f(x)?

OpenStudy (anonymous):

at vertex f'(x)=0 so 2x-8=0 x=4 so y=(4)^2-8(4)-5 = -21 So that;s the minimum, so y is defined on [-21,infinity)

OpenStudy (anonymous):

what is the range then?

OpenStudy (anonymous):

given above, -21 to positive infinity

OpenStudy (anonymous):

what would the x coordinate be?

OpenStudy (anonymous):

is that the )?

OpenStudy (anonymous):

0*

OpenStudy (anonymous):

the x coordinate at y=-21 is 4, but x can be any real number

OpenStudy (anonymous):

so the vertex is x= 4 and y=-21?

OpenStudy (anonymous):

yeah